A moving-coil galvanometer has a 50-turn coil of size 2 cm × 2 cm. It is suspended between the magnetic poles producing a magnetic field of 0.5 T. Find the torque on the coil due to the magnetic field when a current of 20 mA passes through it.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Given θ = 37° d = 10 cm = 0.1 m We know M/B base H = 4π/μ base 0 (d^2 - ℓ^2)^2/2d tan θ = 4π/μ base 0*d^4/2d tan θ [As the magnet is short] = 4π/4π * 10^-7 * (0.1)^3/2 * tan 37° = 0.5 * 0.75 * 1 * 10^-3 * 10^7 = 0.375 * 10^4 * 3.75 * 10^3 A-m^2 T^-1